A 68-kg skier is going down a slope oriented 32° above the horizontal. The area of each ski in contact with the snow is 0.13 m2. Determine the pressure that each ski exerts on the snow.
Answer #1

P = F/A = m*g*cos32°/(2*0.13) = 2173.612505 Pa